OpenStudy (anonymous):

Which factor can you cancel from the numerator and denominator to simplify 7 times x minus 21 all over 18 minus 6 times x.?

OpenStudy (anonymous):

i just want to know how to set it up

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (anonymous):

??? I believe I said this a few moments ago

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

7 from the numerator and 6 from the denominator

OpenStudy (anonymous):

thats the answer? @TURITW

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

okay.. i have two more questions

OpenStudy (anonymous):

Simplify x squared minus 36 all over x squared minus 12 times x plus 36..

OpenStudy (anonymous):

For first question (7x - 21) / (18 - 6x) = 7(x - 3) / (6(3 - x)) = -7(x - 3) / (6(x - 3)) = -7 / 6

OpenStudy (anonymous):